OpenStudy (anonymous):

whats the derivative of 2x - y^2?

OpenStudy (anonymous):

2-2y

OpenStudy (anonymous):

??? how?

OpenStudy (anonymous):

you have no equation here

OpenStudy (anonymous):

dy/dx=2x-y^2

OpenStudy (anonymous):

sorry your right.

OpenStudy (anonymous):

if it is \[2x-y^2=0\] for example then you can find \(y'\) or \(x'\)

OpenStudy (anonymous):

i know its y''=2-... i have no idea how to differentiate y^2 with respect to x

OpenStudy (anonymous):

@Rezz5 we need a better starting point

OpenStudy (anonymous):

deravative of 2x = 2x^(1-1) = 2x^0 and x^0 =1 so 2*1=2 deravative of y^2 = 2*y^(2-1) = 2y^1 =2y

OpenStudy (anonymous):

@Sjano it cant be 2y,

OpenStudy (anonymous):

\[2x-y^2\] is just an expression it does not define y as a function of x, or x as a function of y it is not an equation at all

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

it is not a curve, it is not anything.

OpenStudy (anonymous):

but whats is its derivative?

OpenStudy (anonymous):

would it be y''=2-2y'y?

OpenStudy (anonymous):

it doesn't have one because y is not a function of x, and x is not a function of y you need an EQUATION of some kind is this a problem from a book?

OpenStudy (anonymous):

you are given y'=2x-y^2 i need y''

OpenStudy (anonymous):

ok now you have an equation right? with an equal sign

OpenStudy (anonymous):

i gave typed it in when you told there was no equation above...

OpenStudy (anonymous):

ok now i see it. i didn't see the \(\frac{dy}{dx}=2x-y^2\) so we are thinking that y is some function of x so it is like having \[f'(x)=2x-f(x)^2\] and we want \[f''(x)\] take the derivative with respect to x and get \[f''(x)=2-2f(x)f'(x)\] or \[y''=2-2yy'\]

OpenStudy (anonymous):

now we know \[y'=2x-y^2\] so we can replace the \(y'\) by \(2x-y^2\) above and get \[y''=2-2y(2x-y^2)\]

OpenStudy (anonymous):

so would the derivative of y^3 with respect to x be (3y^2)*y'?

OpenStudy (anonymous):

yea, exactly

OpenStudy (anonymous):

THANK YOU.. i remember know!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now*